could someone please help me with this problem?
Solve the following system of equations by graphing:
y=x^2-4 and y=x-2
???????

Answer by mukhopadhyay(490)

y=x^2-4
=>y=(x-2)(x+2)
Substituting y=x-2 in the above equation we get,
y=y(x+2)
From the above we get, either y = 0 or (x+2) = 1(dividing both sides by y)
x+2=1=>x=-1;
If x=-1, y =x-2=>y=-1-2=-3
If y=0, y=x-2=>0=x-2=>x=2
The answers are: either x=-1 and y=-3; or x=2 and y=0
You can put the solutions in the original equation to verify